half life of a radioactive substance is 20 minute . the time between 20% and 80% decay will be ?

Grade:12th pass

1 Answers

divya bindal
6 years ago
We have, N/N°=1/2^(t/t^1/2)For 20% decay80/100=1/2^(t1/20)Taking log we get.09=0.3t1/20So t1= 6 minFor 80% decay20/100= 1/2^(t2/20)Taking log we get.69=0.3t2/20So t2= 46minSo time between decayt1-t2= 40min